The following data are extracted from the current year’s budget of J.J. Trading.


Machine   shop

Assembly   shop

Finishing   shop

Direct labour hours

20,000

18,000

10,000

Hourly wage rate

RM3.25

RM3.10

RM3.00

Production overhead

RM70,000

RM45,000

RM20,000

 

A customer has requested a quotation for supplying a special piece of equipment and the estimator has worked out the following details:

Components to be purchased

RM216

Materials from store

RM270

Direct labour times:


     Machining

80   hours

     Assembling

40   hours

     Finishing

30   hours

Delivery cost to customer

RM60

 

Administration and marketing costs are covered by adding 10% to cost of manufacture. 12 ½% is added to the total (delivered) cost for profit.

You are required to :

(a) Calculate the cost for supplying this special piece of equipment.

(b) Calculate the selling price quoted to the customer.
Answer #1

Requirement a:

Cost for supplying special piece of Equipment RM1,600

Working notes:

Overhead rates

Machine shop = RM70,000 / 20,000 = RM3.50

Assembly shop = RM45,000 / 18,000 = RM2.50

Finishing shop = RM20,000 / 10,000 = RM2.00

Total Cost

RM
Components to be purchased 216
Materials from Store 270
Direct Labor Cost
Machining (80 x RM3.25) 260
Assembling (40 x RM3.10) 124
Finishing (30 x RM3.00) 90
Overhead Cost
Machining (80 x RM3.50) 280
Assembling (40 x RM2.50) 100
Finishing shop (30 x RM2.00) 60
Total Manufacturing Costs 1,400
Delivery cost 60
Administrative and Marketing cost (RM1,400 x 10%) 140
Total Cost 1,600

Requirement b:

Selling Price RM1,800

Working notes:

Total Cost RM1,600
Add: Profit (RM1,600 x 12.5%) RM200
Selling Price RM1,800

All the best...